How to calculate those problems?

1. The monetary base (H) is $ 1,000; the currency ratio is 25 per cent; the reserve ratio is 10 per cent. What is the dollar amount of deposits?

Not more than $ 3,000

More than $ 3,000 but not more than $ 3,250

More than $ 3,250 but not more than $ 3,500

More than $ 3,500

2. The monetary base (H) is $ 1,000; the currency ratio is 40 per cent; the reserve ratio is 25 per cent. What is the total amount of loans?

More than $ 2250

More than $ 2000, but not more than $ 2250

More than $ 1500, but not more than $ 1750

Not more than $ 1500

3.

The monetary base (H) is $ 1,000; the currency ratio is 25 per cent; the reserve ratio is 10 per cent. What is the dollar amount of reserves?

Not more than $ 250

More than $ 250 but not more than $ 300

More than $ 300 but not more than $ 350

More than $ 300

4. The monetary base (H) is $ 1,000; the currency ratio is 40 per cent; the reserve ratio is 25 per cent. What is the total amount of reserves?

More than $ 225

More than $ 200, but not more than $ 225

More than $ 150, but not more than $ 175

Not more than $ 150
Answer #1

1) Monetary base = C + R

The monetary base (B) is money held by the public in currency and by banks as reserves R

CR = C/D

The currency-deposit ratio (cr) is the amount of currency people hold as a fraction of their demand deposits

So, H = C + R

R = 10% of deposits = 0.10D

So, C = H-R

= C = H-0.10D

Now, CR = C/D

= 0.25 = H-0.10D/D

= 0.25D = 1000 - 0/10D

So, 0.35D = 1000

D = $2,857.14

So, this is not more than $3,000

2) Loan amount is remaining balace after reducing the reserves from the deposits

So, C = H-R

R = 0.25D

C = H-0.25D

CR = C/D

0.40 = 1000 - 0.25D/D

= 0.65D = 1000

D = $1,538.46

So, loans = 1,538.46 - 25% reserves

= Loans = $1,153.85

This is not more than $1,500

3. This calculation is same as question 1.

D = $2,857.14

Reserves = 10%D

= $285.714 (2,857.14 * 10%)

This is more than $250 but less than $300

4. This is same as question 2.

D = $1,538.46

Reserves = 25%D

= $384.615 (1538.46*25%)

This is more than $225